OpenStudy (anonymous):

Find a point on the graph of y= e^(2x) at which the tangent line passes through the origin.

OpenStudy (anonymous):

This was a tough one! First find the derivative. y' = 2e^(2x) The tangent line is y = mx+b or y = 2e^(2x)*x+b, since y' = slope = m The initial value of the tangent line is 0 since it passes through the origin, so b=0. y = 2e^(2x)*x The tangent line crosses with the original graph, so you can set them equal to each other. 2e^(2x)*x = e^(2x) 2x=1 x=1/2 There's your point :) This was tricky, feel free to ask if you don't understand